OpenStudy (marigirl):

How to get from (r-1)/r to 1-(1/r) by simplifying.

OpenStudy (anonymous):

\[\frac{ (r-1) }{ r } = \frac{ r}{ r }-\frac{ 1 }{ r }\]\[\frac{ r }{ r }=1\]Therefore\[\frac{ (r-1 }{ r }=1-\frac{ 1 }{ r }\]

OpenStudy (anonymous):

you can understand that?

OpenStudy (marigirl):

yes! thank you!!

OpenStudy (anonymous):

yw